Sailright Inc. manufactures and sell sailboards. Management believes that the price elasticity of demand is -3.0. Currently, boards are priced at $500 and the quantity demanded is 10000 per year. If the price is increased to $600, how many sailboards will the company be able to sell each year? How much will total revenue change as a result of the price increase?


1
Expert's answer
2022-03-22T15:53:16-0400

Ed="\frac{ change of quantity demanded}{ change in price}"


% change in Quantity


"\frac{Q1-Q}{Q}" ="\frac{Q1-10000}{10000}"


=0.0001Q1-1


% change in price


"\frac{P1-P}{P}" ="\frac{600-500}{500}"

="\frac{100}{500}" =0.2


Ed= "\frac{0.0001Q1-1}{0.2}" ,Where Ed=-3,then


-3="\frac{0.0001Q1-1}{0.2}"


-0.6=0.0001Q1-1


0.4=0.0001Q1

Q1=4000

change in revenue ;

=(P1"\times" Q1)-(P"\times"Q)

=(600"\times"4000)-(500"\times"10000)

=2,400,000-5,000,000

=-2,600,000

therefore by increasing prices the total revenues reduced by $2,600,000